OpenStudy (anonymous):

can you please answer the 7 questions below???

OpenStudy (anonymous):

1. There are 18 boys and 12 girls in a math class.what is the ratio of girls to total students? A.18:30 B.12:30 C.30:12 D.30:18 2.which of the following ratios are equivalent to 6:9? Choose all that apply. A.3:2 B.18/27 C.30:45 D.24/32 3.what is the unit price for one CD if CDs cost 30.00? A. $9.00 B. $8.00 C. $10.00 D. $9.50 4.milo types 35 words per minute.how many words can he type in 7 minutes? A.245 words B.240 words C.250 words D.275 words 5.select the pairs of ratios that form a proportion. A. 3:6 and 5:10 B.4/12 and 6/18 C. 20/5 and 10/3 D. 24:36 and 8:9 6. Solve the proportion 2/3=8/P A. P=14 B p=12 C. P=48 D. P=50 7.which is the better buy? 1. 5 oranges for 1.60 2.7 oranges for 2.45 A.1 B.2 C.They are both equally good values.

OpenStudy (anonymous):

The answers are on yahoo http://answers.yahoo.com/question/index?qid=20121228133718AAn55KG

OpenStudy (anonymous):

oh wow thanks

OpenStudy (anonymous):

My pleasure!! I don't think the one for 3 is the same as the one asked on yahoo but it should atleast give you an idea of how to figure it out:D